Project

General

Profile

12022013

Nightly Observing Plan for MJD 56629 (starting evening of Dec 2, 2013)

Observers

Observer 1 Observer 2 Run Manager Assistant Run Manager SISPI Expert on Call
Rutu Das Marcelle Soares-Santos Kevin Reil Lyndsay Old
ext. 521 ext. 518 ext. 513 or 630 338 9267 ext. 516

CTIO staff (Telescope Operators)

Day Support Night Assistant
Manuel Hernandez Claudio Aguilera

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56629, date 20131202, Site CTIO, UT-offset 3 hr
Current local time 16:51 CLST     UT 19:51     SidTim 19:56
Sunset     20:39 CLST  23:39 UT   SidTim  356 deg  
Eve. -10   21:19 CLST  00:19 UT   SidTim    6 deg  Zeropoint; standards
Eve. -14   21:42 CLST  00:42 UT   SidTim   12 deg  Begin obstac
Eve. -18   22:06 CLST  01:06 UT   SidTim   18 deg  Last chance standards
Midpoint   01:33 CLST  04:33 UT   SidTim   69 deg  
Morn -18   04:59 CLST  07:59 UT   SidTim  121 deg  Begin astronomical twilight
Morn Std   05:23 CLST  08:23 UT   SidTim  127 deg  End obstac; begin standards
Morn -10   05:46 CLST  08:46 UT   SidTim  132 deg  
Sunrise    06:26 CLST  09:26 UT   SidTim  143 deg  
Moon position (deg): RA 252.5, Dec -18.6 Phase: 0.004

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards (preceed with zeros script: script to use:**flush_Vsub.json**).
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/12/03 00:19' Local time = 21:19
SDSSJ0320-0000 ra 03:20:00 dec +00:00:00 secz 1.6 script to use: DES_std_sdssj0320_Yzirgu
SDSSJ0000-0000 ra 00:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0000_Yzirgu

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/12/03 08:23' Local time = 05:23
E2-A ra 04:03:00 dec -44:41:45 secz 1.7 script to use: DES_std_e2a_ugrizY
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj1048_ugrizY
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_ugrizY

Notes from 4pm Meeting

Telescope & Camera
------------------
Telescope is OK. camera is OK. Telescope on white spot.
PTC's in progress now.
System 1 hard drive was replaced today.
Last Night
----------
Smooth observing. On shutter error. WF & SN. SN is all current.
No noise, no hexapod problems, S30 dead on one side, bright source in
other corner. Marco and Juan planning to test it. Need to get the
test-stand fired up to test the electronic card.
DESDM
-----
11/28 is DQ WF is done, by the time we got to WF the noise had gone away
11/30 is running
12/1 to be started soon.
New version of the assessment script is running and Robert has a 
separate DB table through Nov 15th. This is seeing FWHM, background,
clouds. DQ is on the 1st two. There is also an "unknown".
Differs most in Y and z, but less in the others.
There are a problematic types of images: too much extinction to
process it (fail), deep fields are sometimes too good (all cataloged
stars saturated) (good).
SN DQ for 11/28 was applied last night.
Tonight
--------
This afternoon we're getting a PTC, triggered by crate 1 PS replacement.
Weather Looks great. New Moon.
Obstac does Tiling 3 gri, probably.
AOB
---
Add the u-band to one of the SS for tonight.
Do it again on Thursday after they clean the primary.
Have CCD counts to Photodiode counts changed? Ask WW and DJ.